OpenStudy (anonymous):

The hypotenuse of a right triangle is 25 cm long. One leg of the same triangle is 7 cm long. What is the length of the other leg?

OpenStudy (anonymous):

24

OpenStudy (anonymous):

by using pythagoras theorem a^2=c^2-b^2 c=hypotenuse a^2=25^2-7^2 a^2=625-49 a^2=576 a=sqrt576 a=24